OpenStudy (anonymous):

locate the absolute extrema for f(x)=(x+2)^1/3 on [-4,0] i have the min which i got -1.295 and the max 1.295. but i am not 100 percet sure, help please

OpenStudy (anonymous):

if you just need to check your answer, use a graphing calculator. A TI-83/84 is should be handy

OpenStudy (anonymous):

well that it what i used, with the -4 and 0, but i dont get the same answer on paper

OpenStudy (anonymous):

what is your derivative?

OpenStudy (anonymous):

\[1/3(x+2)^-2/3\]

OpenStudy (anonymous):

-2/3 is a power

OpenStudy (anonymous):

(1/3)(x+2)^(-2/3). What did you do next?

OpenStudy (anonymous):

i plugged in the vaules you got into x which were -4 through 0

OpenStudy (anonymous):

?? plugged -4 and 0 in for x of which function? the derivative function or the function itself?

OpenStudy (anonymous):

like this 1/3(-4+2)^(-2/3), i plugged in -4 from [-4, 0], for x

OpenStudy (anonymous):

you can't do that. Or what you didn't isn't helpful. What you were supposed to do were to find the critical number by setting the derivative equal to 0

OpenStudy (anonymous):

what you did*

OpenStudy (anonymous):

so then how would i go with that, i am kinda lost right now

OpenStudy (anonymous):

did you do f'(x) = 0? You already found f'(x)

OpenStudy (anonymous):

yes, but then i confused what would you do to get x alone i mean i know you are suposed to divide it or subtract it, but how with this one

OpenStudy (anonymous):

how would you solve (1/3)(x+2)^(-2/3) = 0 ?

OpenStudy (anonymous):

yes would you subtract 1/3 from both sides

OpenStudy (anonymous):

no

OpenStudy (anonymous):

multiply both sides by 3

OpenStudy (anonymous):

so then you would get 0=(x+2)^(-2/3)

OpenStudy (anonymous):

yes, that is equivalent to 0 = 1/(x+2)^(2/3)

OpenStudy (anonymous):

which has no solution. However, -2 would be the critical number since f'(-2) does not exist. So you have to points to check: f(-4), f(-2), f(0) whichever one is the smallest is the abs min whichever one is the largest is the abs max

OpenStudy (anonymous):

so then would you plug in -4 into the x then

OpenStudy (anonymous):

yes but not in f'(x) but f(x), the original function

OpenStudy (anonymous):

ok thats what i thought

OpenStudy (anonymous):

well thank you very much